Presented below are the production data for the first six months of the year for the...

  1. Presented below are the production data for the first six months of the year for the mixed costs incurred by Vertise Company;

Month

Cost

Units

January

$4,890

4,100

February

4,426

3,520

March

6,730

5,200

April

9,724

8,250

May

5,900

4,900

June

7,336

6,600

Vertise Company uses the high-low method to analyze mixed costs. How would the cost function be stated?

  1. Y = $7,850 + $1.18X
  2. Y = $2,388 + $1.12
  3. Y = $5,298 + $1.19X
  4. Y = $484 +$1.12X
  5. None of the above
Answer #1

Variable cost = (Cost at highest activity - Cost at lowest activity) /(Highest activity-Lowest activity)

= (9724-4426)/(8250-3520) = 1.12

Fixed cost = 9724 - (8250*1.12) = 484

The equation would be

Y = 484 + 1.12X

Option D is the answer